There were 126 people that ordered food. 87 ordered cold food and 92 ordered hot food. How many ordered both cold and hot?
Answer by flame8855(424)

126 - 87 = 39 people who ordered hot food only
126-92 = 34 people who ordered cold food only
126 - 39-34= 53 people ordered both